OpenStudy (anonymous):

How do I find the distance on the coordinate system from the point (-3,4) and the point (8,-7)???

hero (hero):

Are you familiar with pythagorean theorem?

OpenStudy (anonymous):

sort of..

hero (hero):

a^2 + b^2 = c^2

hero (hero):

 If you know the lengths of sides a and b, then you can use that formula to find the length of side c.

hero (hero):

In this particular case, you have points (-3,4) and (8,-7) which you can plot and make a triangle: 

hero (hero):

Notice that if you're given points (a,b) and (c,d), then you can figure out a third point (a,d) to create a right triangle along with the lengths of sides a and b. Furthermore, you can determine the lengths of the sides.

hero (hero):

Now you can use a^2 + b^2 = c^2 (pythagorean theorem) to figure out the length of the third side which would give you the distance between the two given points.